OpenStudy (anonymous):

Please help derives the following f (x) = e ^ x [(x-9) / (x-7)] ^ 1 / 2

OpenStudy (amistre64):

is this a product of "e" and "x-9"/"x-7"?

OpenStudy (amistre64):

\[f (x) = (e ^ x)\left(\cfrac{(x-9)}{(x-7)}\right) ^ {1 / 2}\]

OpenStudy (anonymous):

Yuk...

OpenStudy (amistre64):

[rl]' = r'l + rl' ; my right and left get transposed, but i think it looks better that way :) r = e^x ; r'=e^x l=\(\left(\cfrac{x-9}{x-7}\right)^{^1/_2}\) ; l' = a bit complicated :)

OpenStudy (anonymous):

x-7 = u^2

OpenStudy (anonymous):

i did this super fast so could be wrong, but i get: e^x * (x - 7)^(-3/2) * (x - 8)^2 / (x - 9)^(1/2)

OpenStudy (amistre64):

[t b^(-1)] = t b' + t' b t = x-9; t' = 1 b = (x-7)^-1 ; b'= (-x+7)^-2 \(\cfrac{x-9}{(7-x)^2}\) + \(\cfrac{1}{x-7}\) looks to the the derivative of the fractional part; b ut then incorporate the sqrt into it ... ugh

OpenStudy (amistre64):

i got b' mixed up :)

OpenStudy (anonymous):

what kind of sadist is asking this question?

OpenStudy (amistre64):

yeah lol. practice problems are spose to build confidence :)

OpenStudy (anonymous):

e^(u^2+7)(u^2-2)^(1/2)